OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6. 1 4 5 25

OpenStudy (anonymous):

3(x + 1) = -2(x - 1) + 6. 3x+ 3= -2x +2 +6 3x+2x = 8-3 5x= 5 x=?

OpenStudy (anonymous):

1 ?

OpenStudy (ipwnbunnies):

Yes